OpenStudy (anonymous):

solve the equation 9^x = 34

OpenStudy (anonymous):

You would use the natural log to find the answer. ln(9^x) = ln(34) ==> x * ln(9) = ln(34) ==> x = ln(34)/ln(9)

OpenStudy (danjs):

\[\large \log_{b}b^x = x \] is a property

OpenStudy (anonymous):

so confused

OpenStudy (danjs):

you can do it zeldas way using the power property , bringing the x down in front using the property i posted , take the log base 9 of both sides, and then you change the base formula to get the same thing

OpenStudy (anonymous):

\[\huge b^x=A\iff x=\frac{\ln(A)}{\ln(b)}\] and a calculator

OpenStudy (anonymous):

in your case \(A=34,b=9\)

OpenStudy (anonymous):

oh so 1.605?

OpenStudy (anonymous):

idk i didnt do it want to check?

OpenStudy (anonymous):

its right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok wait

OpenStudy (anonymous):

\[16^x = \left(\begin{matrix}1 \\ 4\end{matrix}\right)^{x+3}\]

OpenStudy (anonymous):

how do you do this

OpenStudy (anonymous):

is that sposed to be a fraction?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[16=2^4\]right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

next job is to write \[\frac{1}{4}\] as a power of \(2\) how can you do that?